OpenStudy (anonymous):

(10x+y)(10x-y) simply...:(

OpenStudy (anonymous):

Pls gave me a Replay ♥

hartnn (hartnn):

\(\huge (A+B)(A-B)=A^2-B^2\) use this. put A=10x B=y and tell me what u get.

OpenStudy (anonymous):

okay..Thnx..♥

OpenStudy (amistre64):

i tend to have to wonder what "simplify" means in this context. But i agree that hartnns approach is correct.

Parth (parthkohli):

\[(\phi + \psi)(\phi -\psi) = \phi^2 - \psi^2 \ \]

OpenStudy (amistre64):

lol, but what is it in celtic runes?

Parth (parthkohli):

"Replay" "simply" These words don't make sense to me.

hartnn (hartnn):

i makes sense to me he meant reply simplify

OpenStudy (amistre64):

\[(☺-☻)(☺+☻)=☺^2-☻^2\]

OpenStudy (amistre64):

that might look better larger \[\Huge(☺-☻)(☺+☻)=☺^2-☻^2\]

hartnn (hartnn):

\(\Huge \color{red}{(♥-\clubsuit)(♥+\clubsuit)=♥^2-\clubsuit^2}\) my try :)

OpenStudy (amistre64):

lol, dont make me rainbow it up ;)

hartnn (hartnn):

lol!

Parth (parthkohli):

Whom to give a medal to?

OpenStudy (amistre64):

eenie meenie minee mo ....

OpenStudy (anonymous):

Sry...Guys..For the Mistake..♥